OpenStudy (anonymous):

PLEASE HELP I WILL MEDAL AND FAN The sum of two consecutive even integers is 226. What is the larger integer? 114 116 112 118

OpenStudy (anonymous):

x + x + 2 = 226 Or, 2x = 224 Now can you solve it for x

OpenStudy (mathstudent55):

We can solve this with an equation. We start by choosing a variable. Let x be the smaller number.

OpenStudy (mathstudent55):

The next larger even number is 2 larger than x, that makes the larger number x + 2.

OpenStudy (anonymous):

224/2 made it easier

OpenStudy (mathstudent55):

Since we are told what the sum of the two consecutive even numbers is, we use our expressions with x and we follow what the problem tells us. The sum of the two consecutive even numbers is x + x + 2. We are told the sum is equal to 226, so that gives us our equation: \(x + x + 2 = 226\)

OpenStudy (anonymous):

112

OpenStudy (anonymous):

but thats not answer

OpenStudy (anonymous):

wait would it be 110 and 112?

OpenStudy (mathstudent55):

Now we add the like terms on the left side: \(2x + 2 = 226\) We subtract 2 from both sides: \(2x = 224\) The next step is to divide both sides by 2: \(x = 112\) Now we need to go back to our use of the variable x.

OpenStudy (mathstudent55):

We said above that x is the SMALLER of the two numbers. The problem is asking for the larger of the two numbers. The larger number is x + 2, so 112 + 2 = 114. The answer is: the larger number is 114.

OpenStudy (anonymous):

would it be 110 and 112? no.. 110 and 108???

OpenStudy (anonymous):

nope...

OpenStudy (mathstudent55):

110 + 112 = 222 110 + 108 = 218 Neither of the above add to 226, so they cannot be correct.

OpenStudy (anonymous):

its 226 not 228

OpenStudy (anonymous):

i got 114

OpenStudy (mathstudent55):

Just because you solved the equation for x, it does not mean you have finished. You need to see what x represents and make sure you answer the question that is being asked.

OpenStudy (anonymous):

226/2 is not 114

OpenStudy (mathstudent55):

Follow my responses above to understand how to get the answer.

OpenStudy (anonymous):

wait the answer is 114!?

OpenStudy (mathstudent55):

Yes.

OpenStudy (freckles):

yes they were saying x+2>x for positive number x

OpenStudy (freckles):

and you found the x so you could plug that x into x+2 to find the large of the two

OpenStudy (anonymous):

lol yeah ???? i did not know how to explain it